OpenStudy (anonymous):

verify that the x-vales are solutions of the equation 2cos^2 4x-1=0 (a) x= pi/16 (v) x=3pi/16

OpenStudy (lgbasallote):

try substituting

OpenStudy (anonymous):

yeah idk how to solve after that xD

OpenStudy (anonymous):

2cos^2 4 (pi/16)-1=0

OpenStudy (anonymous):

what does pi/16 equal?

OpenStudy (lgbasallote):

you multily it to 4 what do you get

OpenStudy (lgbasallote):

\[\frac{\pi}{16} \times 4\]

OpenStudy (anonymous):

4π/16

OpenStudy (anonymous):

π/4

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

and what's cos pi/4?

OpenStudy (lgbasallote):

 you are familiar with the 45-45-90 triangle right?